OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (3, 1) and (2, 9)? y = x +

OpenStudy (anonymous):

@princeharryyy

OpenStudy (shamim):

U know slope of ur line is m=(y1-y2)/(x1-x2)=?

OpenStudy (princeharryyy):

y = -8x +c 1 = -24 +c y=-8x + 25